OpenStudy (anonymous):

-10-3/7x=5 solve show work

OpenStudy (anonymous):

-13/35

OpenStudy (anonymous):

i really need to see work

OpenStudy (anonymous):

not jus an answer

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

come on is that correct ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok then you gotta solve it and correct me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

plz right now

OpenStudy (anonymous):

multiply both sides by 7x therefore, you get -10-3=35x -13=35x and x = -35/13

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

almost same

OpenStudy (anonymous):

Are you an instructor?

OpenStudy (anonymous):

use this link, and it will give you the steps to solve your problem http://www.wolframalpha.com/input/?i=-10-3%2F7x%3D5+solve

OpenStudy (anonymous):

wolfram dont always give steps and no krish

OpenStudy (anonymous):

Is wolfram free to use kinda thing?

OpenStudy (anonymous):

It gives steps, just go to the site and it's very helpful